Vijay bought 11 tickets for a waterslide. The tickets were $6 for adults and $4 for children. Vijay paid $60 for the tickets. For how many adults and how many children did he buy tickets?
Answer by mananth(12269)

Adult x tickets
child y tickets

1 x + 1 y = 11 .............1
6 x + 4 y = 60 .............2
Eliminate y
multiply (1)by -4
Multiply (2) by 1
-4 x -4 y = -44
6 x 4 y = 60
Add the two equations
2 x = 16
/ 2
x = 8
plug value of x in (1)
1 x + 1 y = 11
8 + 1 y = 11
1 y = 11 -8
1 y = 3
y = 3
Adult 8 tickets
child 3 tickets
m.ananth@hotmail.ca